OpenStudy (anonymous):

4i/2+3i divide and express the result in standard form.

OpenStudy (anonymous):

rationalise the denominator.

OpenStudy (tkhunny):

4i/2 + 3i = 2i + 3i = 5i Did you mean 4i/(2+3i)? That is not what you have written.

OpenStudy (dumbcow):

multiply top/bottom by conjugate \[\frac{4i(2-3i)}{(2+3i)(2-3i)}\]